Oslo Company prepared the following contribution format income statement based on a sales volume of 1,000 units (the relevant range of production is 500 units to 1,500 units):

Sales $ 25,000
Variable expenses 17,500
Contribution margin 7,500
Fixed expenses 4,200
Net operating income $ 3,300

1. What is the variable expense ratio?

2. What is the contribution margin per unit? (Round your answer to 2 decimal places.)

3. What is the contribution margin ratio?

4. What is the variable expense ratio?

5.If sales increase to 1,001 units, what would be the increase in net operating income? (Round your answer to 2 decimal places.)

6.If sales decline to 900 units, what would be the net operating income?

7.If the selling price increases by $2 per unit and the sales volume decreases by 100 units, what would be the net operating income?

8.If the variable cost per unit increases by $1, spending on advertising increases by $1,150, and unit sales increase by 130 units, what would be the net operating income?

9.What is the break-even point in unit sales?

10.

What is the break-even point in dollar sales?

11.How many units must be sold to achieve a target profit of $4,500?

12.What is the margin of safety in dollars? What is the margin of safety in percentage

13. What is the degree of operating leverage? (Round your answer to 2 decimal places.)

13.

9.

10.
Answer #1

1. Variable expense ratio = (17.5 / 25.00) x 100 = 70%

2. Contribution margin per unit = 25.00 - 17.50 =$ 7.50

3. Contribution margin ratio = (7.50 / 25.00) x 100 = 30%

4. Variable expense ratio = (17.5 / 25.00) x 100 = 70%

5.If sales increase to 1,001 units, the increase in net operating income will be $ 7.50

6.If sales decline to 900 units, the net operating income = 900 x 7.50 - 4200 = $ 2,550

7.If the selling price increases by $2 per unit and the sales volume decreases by 100 units, the net operating income = 1100 x 9.50 - 4200 = $ 6,250

8.If the variable cost per unit increases by $1, spending on advertising increases by $1,150, and unit sales increase by 130 units, the net operating income = 1130 x 6.50 - 5350 = $ 1,995

9. The break-even point in unit sales = 4200 / 7.50 = 560

10. The break-even point in dollar sales = 560 x 25.00 = $14,000

11.How many units must be sold to achieve a target profit of $4,500 = (4200 + 4500) / 7.50 = 1160

12. The margin of safety in dollars = 25,000 - 14,000 = $ 11,000

The margin of safety in percentage = (11000/25000) x 100 = 44%

13. Degree of operating leverage = 7500 / 3300 = 2.27